OpenStudy (anonymous):

A 25.0g sample of magnesium at 22.0ºC absorbs 356.3 J of heat. The temperature of the magnesium rises to 36.5ºC. What is the specific heat of the magnesium?

OpenStudy (anonymous):

q=cm(T2-T1) 356.3J = c(25g)(14.5°C) c=356.3/(25x14.5) c=0.98 J/g°C But, SI units should be J/kgK. So it should be ~3.57 J/kgK if my conversion is correct.